# Unifying Collisional Models and the Monte Carlo Metropolis Method: Algorithms for Dynamics of Open Quantum Systems

### v1

@@ IV.1 Collisional model: Single time step evolution
$\begin{split}\rho_{S}(t_{n+1})&=\rho_{S}(t_{n})+\frac{(g\Delta t)^{2}}{Z_{a}}\sum_{i>j}\bigg{[}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}e^{-\beta\epsilon_{0}}\\ &+A_{i,j}^{\dagger}\rho_{S}(t_{n})A_{i,j}e^{-\beta\epsilon_{\alpha_{i,j}}}-\frac{e^{-\beta\epsilon_{0}}}{2}\Big{\{}\rho_{S}(t_{n}),\left|i\right\rangle\left\langle i\right|\Big{\}}\\ &\qquad-\frac{e^{-\beta\epsilon_{\alpha_{i,j}}}}{2}\Big{\{}\rho_{S}(t_{n}),\left|j\right\rangle\left\langle j\right|\Big{\}}\bigg{]},\end{split}$ (17)
where $\left\{A,B\right\}\equiv AB+BA$ is the standard anticommutator.
@@
In order to compare directly with the dynamics generated by the Metropolis algorithm, as detailed in the next section, we must make one additional assumption about the system density matrix, $\rho_{S}(t_{n})$. As discussed in Section III the Metropolis algorithm constructs a Markov chain of transitions between the system’s energy eigenstates. Thus, the density matrix at any given time step of the algorithm will be a weighted average of the energy eigenstates. For an even comparison, we therefore assume that the system density matrix in the collisional model is also diagonal in the energy eigenbasis, $\rho_{S}(t_{n})=\sum_{l}p_{l}(t_{n})\left|l\right\rangle\left\langle l\right|$.
Plugging in this decomposition for the density matrix, Eq. (17) simplifies to,
$\begin{split}&\rho_{S}(t_{n+1})=\rho_{S}(t_{n})+\frac{(g\Delta t)^{2}}{Z_{a}}\sum_{i>j}\Big{[}\Big{(}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}\\ &-p_{i}(t_{n})\left|i\right\rangle\left\langle i\right|\Big{)}+\left(A_{i,j}^{\dagger}\rho_{S}(t_{n})A_{i,j}-p_{j}(t_{n})\left|j\right\rangle\left\langle j\right|\right)e^{-\beta\epsilon_{\alpha_{i,j}}}\Big{]}\end{split}$ (18)
Where, without loss of generality, we have also set the ground state energy of the bath at zero $\epsilon_{0}=0$.
@@ IV.2 Metropolis algorithm: Single time step evolution
Accounting for the accept and reject possibilities, the average system state after a single time step is,
$\begin{split}\rho_{S}(t_{n+1})&=\frac{1}{L}\sum_{i\neq j}\Big{[}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}e^{-\beta\omega_{i,j}}\\ &+\left(1-e^{-\beta\omega_{i,j}}\right)P_{i}\rho_{S}(t_{n})P_{i}\Big{]}\end{split}$ (21)
Note that the projectors $P_{i}\equiv\left|i\right\rangle\left\langle i\right|$ in the second term of Eq. (21) are necessary in order to ensure that the rejection probability $1-e^{-\beta\omega_{i,j}}$ for the jump $\left|i\right\rangle\rightarrow\left|j\right\rangle$ only contributes to the summation when $\rho_{S}(t_{n})$ has some population in eigenstate $\left|i\right\rangle$. Equation (21) simplifies to,
$\begin{split}\rho_{S}(t_{n+1})=\rho_{S}(t_{n})+&\frac{1}{L}\sum_{i\neq j}\Big{[}\Big{(}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}\\ &-P_{i}\rho_{S}(t_{n})P_{i}\Big{)}e^{-\beta\omega_{i,j}}\Big{]}\end{split}$ (22)
@@
$\begin{split}\rho_{S}(t_{n+1})&=\rho_{S}(t_{n})+\frac{1}{L}\sum_{i>j}\Big{[}\Big{(}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}\\ &-P_{i}\rho_{S}(t_{n})P_{i}\Big{)}\Big{]}+\frac{1}{L}\sum_{i<j}\Big{[}\Big{(}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}\\ &\qquad-P_{i}\rho_{S}(t_{n})P_{i}\Big{)}e^{-\beta(E_{j}-E_{i})}\Big{]}\end{split}$ (23)
Noting that $A_{i,j}=A_{j,i}^{\dagger}$ we can swap the indicies in the second summation of Eq. (23) rewrite it as a single summation,
$\begin{split}\rho_{S}(t_{n+1})&=\rho_{S}(t_{n})+\frac{1}{L}\sum_{i>j}\Big{[}\left(A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}-P_{i}\rho_{S}(t_{n})P_{i}\right)\\ &+\left(A_{i,j}^{\dagger}\rho_{S}(t_{n})A_{i,j}-P_{j}\rho_{S}(t_{n})P_{j}\right)e^{-\beta(E_{i}-E_{j})}\Big{]}\end{split}$ (24)
Under the Metropolis scheme, each experimental run consists of a time series of eigenstates. The density matrix at a particular time step, $\rho_{S}(t_{n})$ is then constructed by averaging over many experimental runs. With this in mind, we can in general represent the density matrix as a weighted average of the system eigenstates,
$\rho_{S}(t_{n})=\sum_{l}p_{l}(t_{n})\left|E_{l}\right\rangle\left\langle E_{l}\right|.$ (25)
Plugging Eq. (25) into Eq. (24) yields,
$\begin{split}\rho_{S}(t_{n+1})&=\rho_{S}(t_{n})+\frac{1}{L}\sum_{i>j}\Big{[}\left(A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}-p_{i}(t_{n})\left|i\right\rangle\left\langle i\right|\right)\\ &+\left(A_{i,j}^{\dagger}\rho_{S}(t_{n})A_{i,j}-p_{j}(t_{n})\left|j\right\rangle\left\langle j\right|\right)e^{-\beta(E_{i}-E_{j})}\Big{]}\end{split}$ (26)
Recalling that, in the collisional model, we structured the bath energies such that $E_{i}-E_{j}=\epsilon_{\alpha_{i,j}}$ we can compare Eqs. (26) and (18) and see that they are identical when the condition $1/L=(g\Delta t)^{2}/Z_{a}$ is satisfied.
@@ IV.3 Conditions of equivalence
In this subsection we will consider in detail the conditions necessary to fulfill the equivalence demonstrated in the previous two subsections. Analytically, we have shown that the discrete time evolution of the density matrix generated by the collisional model, Eq. (18), and the Metropolis algorithm, Eq. (26) have exactly the same structure. However, we note that the truncated expansion in Eq. (16) used to derive the time-evolved collisional model density matrix is only accurate under the condition that $g\Delta t\ll 1$. Furthermore, the collisional model bath ancillae must be structured in a thermal state that fulfills the transition energy matching condition. Finally, as noted previously, the exact equivalence between Eqs. (26) and (18) requires $1/L=(g\Delta t)^{2}/Z_{a}$.
To compare the validity of this equivalence, in Fig. 6 we plot the trace distance between the density matrix generated by Eq. (11) for the collisional model and the density matrix arising from averaging over 100,000 runs of the Metropolis scheme for the one-dimensional XXZ model. Specifically, we use Eq. (11) for the collisional model, rather than Eq. (18), as we wish to test the regimes in which the approximations that went into the derivation of Eq. (18) are valid. We see that the trace distance initially increases, before dropping back to zero, as both models ultimately result in the thermal state density matrix. We also see that the amount by which the trace distance increases depends on the system size, with shorter length chains having a significantly lower maximum trace distance.
Figure 6: Trace distance between the density matrix arising from the collisional model and the density matrix arising from the Metropolis algorithm as a function of model time steps, $n$, for a one-dimensional XXZ chain of length $N=2$ (red, solid), $N=3$ (blue, long dashed), and $N=4$ (green, short dashed). Parameters are $J=h=\Delta=1$ and $\beta=2$. For the collisional model we have fixed the interaction parameter $g\Delta t=\sqrt{Z_{a}/L}$ and for the Metropolis algorithm we have constructed the density matrix from averaging over 100,000 runs.
In our comparison, we have fixed the collisional model interaction parameter using the condition $(g\Delta t)^{2}=Z_{a}/L$. Thus, the only source of the deviation between the density matrices for both models comes from the fact that the truncated expansion of the time evolution operator does not fully capture the dynamics of the collisional model. In order to satisfy both the conditions $g\Delta t\ll 1$ as well as $1/L=(g\Delta t)^{2}/Z_{a}$ we see that we want the ratio of $Z_{a}/L$ to be as small as possible. In Fig. 7 we plot this ratio as a function of chain length for different temperatures. At low temperatures, the ratio remains relatively flat with system size, while at high temperatures it grows exponentially with system size.
@@
Thus, as the system size grows, the ratio $Z_{a}/L$, and thus also $g\Delta t$ grow, rendering the approximate expansion for the time evolution operator increasingly inaccurate and leading to the different dynamics between the models observed in Fig. 6.
@@
On the other hand, in the zero temperature limit, the only terms that contribute to $Z_{a}$ are those where $\epsilon_{\alpha_{i,j}}=0$. This occurs in the case of the ground state energy of the bath and for any $\epsilon_{\alpha_{i,j}}$ corresponding to degenerate pairs of system energy eigenvalues. As the number of degenerate energy states depends on the system size and model parameters, this leads to the non-monotonic behavior of $Z_{a}/L$ observed at low temperatures in Fig. 7. In general, as long as the number of pairs of degenerate eigenvalues grows slower than the total number of eigenvalues, the ratio $Z_{a}/L$ will remain small at low temperatures.
@@
Figure 7: Ratio of the partition function for the collisional model bath ancillae, $Z_{a}$, to the number of possible eigenstate transitions at each time step in the Metropolis algorithm, $L$, for the one-dimensional XXZ model as a function of chain length $N$ at inverse temperature $\beta=200$ (blue, long dashed), $\beta=20$ (red, dotted), $\beta=2$ (green, dot-dashed), $\beta=0.2$ (orange, short dashed), and $\beta=0$ (black, solid). Parameters are $J=h=\Delta=1$.
@@ V Concluding Remarks
In this work we have verified that both the collisional model framework and the Metropolis algorithm lead to thermalization of a many-body system when, in the case of the collisional model, the spectrum of the bath ancillae corresponds to each of the energy eigenstate transitions in the system. We have then demonstrated analytically that not only do both schemes produce thermalization, but that the time-dependent dynamics generated by both models are exactly equivalent when a condition relating the collisional model interaction strength and the ratio of bath ancilla partition function to number of possible energy eigenstate transitions is fulfilled. We have shown that this condition is more easily fulfilled at small system sizes.
@@
Despite the similarities discussed here, there remain some important distinctions between the collisional model and the Metropolis algorithm. By microscopically modeling the system-bath interaction, the collisional model is significantly more general and can be applied with arbitrarily structured bath ancillae to study non-thermal steady states of open system dynamics. Our quantum implementation of the Metropolis scheme also assumes that the system transitions only between individual eigenstates. The collisional model, on the other hand, can model dynamics where the system density matrix has non-zero coherences.
@@
Computationally, the Metropolis algorithm is significantly cheaper as it simply tracks individual eigenstate transitions and needs to know only about the Hilbert space of the system. The collisional model must take into account the much larger joint system-bath Hilbert space and perform repeated partial traces after each system-bath interaction. For example, consider a system whose Hilbert space dimension is $d$. For modeling thermalizing dynamics using a bath whose spectrum has an excited state corresponding to each transition in the system, as we have done in our demonstration of thermalization in the XXZ model, the Hilbert space of each bath ancilla must be of dimension $\binom{d}{2}+1$. Thus the total Hilbert space that the collisional model must operate in is of dimension $\mathcal{O}(d^{3})$. Even in the most optimal scenario, when the spectra of the bath ancillae are identical to the spectrum of the system, the joint system bath Hilbert space will be of dimension $\mathcal{O}(d^{2})$. We must also keep in mind that the Metropolis scheme does incur an additional cost from the fact that the dynamics must be averaged over many repeated runs in order to construct the time-dependent density matrix, while the the collisional model directly models the density matrix evolution. However, each Metropolis experimental run still consists of only a series of products of $d$-dimensional operators.
@@
These results open several avenues for potential future work. While we have shown an equivalence between the collisional model and Metropolis approaches in thermalizing dynamics, it may be possible to generalize it to generic open system dynamics by modifying the distribution that the accept/reject probabilities are drawn in the Monte Carlo approach. Collisional models have seen extensive use in studying non-Markovian open system dynamics. It would be interesting to see if a similar equivalence could be found between non-Markovian collisional models and non-Markovian Monte Carlo schemes for quantum evolution [44, 45, 46].

### v2

@@ IV.1 Collisional model: Single time step evolution
$\begin{split}\rho_{S}(t_{n+1})&=\rho_{S}(t_{n})+\frac{(g\Delta t)^{2}}{Z_{a}}\sum_{i>j}\bigg{[}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}\\ &+A_{i,j}^{\dagger}\rho_{S}(t_{n})A_{i,j}e^{-\beta\epsilon_{\alpha_{i,j}}}-\frac{1}{2}\Big{\{}\rho_{S}(t_{n}),\left|i\right\rangle\left\langle i\right|\Big{\}}\\ &\qquad-\frac{e^{-\beta\epsilon_{\alpha_{i,j}}}}{2}\Big{\{}\rho_{S}(t_{n}),\left|j\right\rangle\left\langle j\right|\Big{\}}\bigg{]},\end{split}$ (17)
where $\left\{A,B\right\}\equiv AB+BA$ is the standard anticommutator. Note that, without loss of generality, we have also set the ground state energy of the bath at zero $\epsilon_{0}=0$.
@@ IV.2 Metropolis algorithm: Single time step evolution
In the typical Metropolis scheme, if a transition is rejected, the state of the system is left unchanged. However, in order to properly incorporate decoherence we modify this procedure to instead apply the map $\mathcal{L}_{i}[\rho]=\frac{1}{2}(\left|i\right\rangle\left\langle i\right|\rho+\rho\left|i\right\rangle\left\langle i\right|)$. Thus the average system state after a rejection is,
$\frac{1}{L}\sum_{i\neq j}\mathcal{L}_{i}[\rho_{S}(t_{n})]=\frac{1}{2L}\sum_{i\neq j}(\left|i\right\rangle\left\langle i\right|\rho_{S}(t_{n})+\rho_{S}(t_{n})\left|i\right\rangle\left\langle i\right|)$ (20)
Accounting for both the accept and reject possibilities, the average system state after a single time step is,
$\begin{split}\rho_{S}(t_{n+1})&=\frac{1}{L}\sum_{i\neq j}\Big{[}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}e^{-\beta\omega_{i,j}}\\ &+\left(1-e^{-\beta\omega_{i,j}}\right)\frac{1}{2}(\left|i\right\rangle\left\langle i\right|\rho_{S}(t_{n})+\rho_{S}(t_{n})\left|i\right\rangle\left\langle i\right|)\Big{]}\end{split}$ (21)
Equation (21) simplifies to,
$\begin{split}\rho_{S}(t_{n+1})=\rho_{S}(t_{n})+&\frac{1}{L}\sum_{i\neq j}\Big{[}\Big{(}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}\\ &-\frac{1}{2}\Big{\{}\rho_{S}(t_{n}),\left|i\right\rangle\left\langle i\right|\Big{\}}\Big{)}e^{-\beta\omega_{i,j}}\Big{]}\end{split}$ (22)
@@
$\begin{split}\rho_{S}(t_{n+1})&=\rho_{S}(t_{n})+\frac{1}{L}\sum_{i>j}\Big{[}\Big{(}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}\\ &-\frac{1}{2}\Big{\{}\rho_{S}(t_{n}),\left|i\right\rangle\left\langle i\right|\Big{\}}\Big{]}+\frac{1}{L}\sum_{i<j}\Big{[}\Big{(}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}\\ &\qquad-\frac{1}{2}\Big{\{}\rho_{S}(t_{n}),\left|i\right\rangle\left\langle i\right|\Big{\}}\Big{)}e^{-\beta(E_{j}-E_{i})}\Big{]}\end{split}$ (23)
Noting that $A_{i,j}=A_{j,i}^{\dagger}$ we can swap the indices in the second summation of Eq. (23) rewrite it as a single summation,
$\begin{split}\rho_{S}(t_{n+1})&=\rho_{S}(t_{n})+\frac{1}{L}\sum_{i>j}\bigg{[}A_{i,j}\rho_{S}(t_{n})A_{i,j}^{\dagger}\\ &+A_{i,j}^{\dagger}\rho_{S}(t_{n})A_{i,j}e^{-\beta(E_{i}-E_{j})}-\frac{1}{2}\Big{\{}\rho_{S}(t_{n}),\left|i\right\rangle\left\langle i\right|\Big{\}}\\ &\qquad-\frac{e^{-\beta(E_{i}-E_{j})}}{2}\Big{\{}\rho_{S}(t_{n}),\left|j\right\rangle\left\langle j\right|\Big{\}}\bigg{]}.\end{split}$ (24)
Recalling that, in the collisional model, we structured the bath energies such that $E_{i}-E_{j}=\epsilon_{\alpha_{i,j}}$ we can compare Eqs. (24) and (17) and see that they are identical when the condition $1/L=(g\Delta t)^{2}/Z_{a}$ is satisfied. Implicit in this equivalence is the premise that the duration of the time step, $t_{n+1}-t_{n}$, is the same for both the Metropolis algorithm and the collisional model. As the Metropolis algorithm is a purely phenomenological model, the time step duration is simply a model parameter. Thus we are free to choose it be identical to the physically-motivated time interval used in the collisional model, as discussed in Section II.
@@ IV.3 Conditions of equivalence
In this subsection we will consider in detail the conditions necessary to fulfill the equivalence demonstrated in the previous two subsections. Analytically, we have shown that the discrete time evolution of the density matrix generated by the collisional model, Eq. (17), and the Metropolis algorithm, Eq. (24) have exactly the same structure. Notably, this equivalence does not explicitly depend on a particular system size or choice of model Hamiltonian. However, we note that the truncated expansion in Eq. (16) used to derive the time-evolved collisional model density matrix is only accurate under the condition that $g\Delta t\ll 1$. Furthermore, the collisional model bath ancillae must be structured in a thermal state that fulfills the transition energy matching condition. Finally, as mentioned previously, the exact equivalence between Eqs. (24) and (17) requires $1/L=(g\Delta t)^{2}/Z_{a}$.
In order to satisfy both the conditions $g\Delta t\ll 1$ as well as $1/L=(g\Delta t)^{2}/Z_{a}$ we see that we want the ratio of $Z_{a}/L$ to be as small as possible. In Fig. 6 we plot this ratio as a function of chain length for different temperatures. At low temperatures, the ratio remains relatively flat with system size, while at high temperatures it grows exponentially with system size.
@@
Here we see that as the system size grows, the ratio $Z_{a}/L$, and thus also $g\Delta t$ grow, rendering the approximate expansion for the time evolution operator increasingly inaccurate and leading to the different dynamics between the models observed in Fig. 7.
@@
On the other hand, in the zero temperature limit, the only terms that contribute to $Z_{a}$ are those where $\epsilon_{\alpha_{i,j}}=0$. This occurs in the case of the ground state energy of the bath and for any $\epsilon_{\alpha_{i,j}}$ corresponding to degenerate pairs of system energy eigenvalues. As the number of degenerate energy states depends on the system size and model parameters, this leads to the non-monotonic behavior of $Z_{a}/L$ observed at low temperatures in Fig. 6. In the optimal scenario, where there are no degenerate eigenvalues, the ratio becomes,
$\lim_{\beta\rightarrow\infty}\frac{Z_{a}}{L}=\frac{1}{2^{N}-1}$ (26)
In general, as long as the number of pairs of degenerate eigenvalues grows slower than the total number of eigenvalues, the ratio $Z_{a}/L$ will remain small at low temperatures.
@@
Figure 6: Ratio of the partition function for the collisional model bath ancillae, $Z_{a}$, to the number of possible eigenstate transitions at each time step in the Metropolis algorithm, $L$, for the one-dimensional XXZ model as a function of chain length $N$ at inverse temperature $\beta=200$ (blue, long dashed), $\beta=20$ (pink, dotted), $\beta=2$ (green, dot-dashed), $\beta=0.2$ (orange, short dashed), and $\beta\rightarrow 0$ (black, solid). The optimal zero temperature lower bound for the case of a non-degenerate spectrum is included for comparison (red, solid). Parameters are $J=h=\Delta=1$.
To compare the validity of the equivalence, in Fig. 7 we plot the trace distance between the density matrix generated by Eq. (11) for the collisional model and the density matrix arising from averaging over $10^{6}$ runs of the Metropolis scheme for the one-dimensional XXZ model. Specifically, we use Eq. (11) for the collisional model, rather than Eq. (17), as we wish to test the regimes in which the approximations that went into the derivation of Eq. (17) are valid. In our comparison, we fix the collisional model interaction parameter using the condition $(g\Delta t)^{2}=Z_{a}/L$. Thus, the only source of the deviation between the density matrices for both models comes from the fact that the truncated expansion of the time evolution operator does not fully capture the dynamics of the collisional model. We see that the trace distance first increases (initially it is zero, as both methods start from the same initial state), before decreasing, as both models ultimately result in the thermal state density matrix.
We also see that the amount by which the trace distance increases depends on the system size. The increase is the largest for a system size of $N=3$, but as system size is increased further to $N=4$ the initial increase in trace distance falls back to the level of the $N=2$ case. This behavior is partially explained by Fig. 6 where we see that, for an inverse temperature of $\beta=20$, the ratio $Z_{a}/L$ decreases between $N=2$ and $N=4$, thus rendering the second order approximation in Eq. (13) increasingly more accurate. The larger trace distance for $N=3$, despite similar values of $Z_{a}/L$, occurs due to the fact that the trace distance is also affected by the choice of initial state, which impacts the rank of the matrix $\rho_{\mathrm{M}}-\rho_{\mathrm{CM}}$ at subsequent time steps. The choice of an initial state that is a pure state composed of an even superposition of all energy eigenstates produces the maximum rank for $\rho_{\mathrm{M}}-\rho_{\mathrm{CM}}$ in the initial dynamics. However, even in this case, we see that the contribution to the trace distance from $Z_{a}/L$ is much more significant, as evidenced by the subsequent decrease in trace distance for $N=4$, which has a lower value of $Z_{a}/L$.
Figure 7: Trace distance between the density matrix arising from the collisional model and the density matrix arising from the Metropolis algorithm as a function of model time steps, $n$, for a one-dimensional XXZ chain of length $N=4$ (red, solid), $N=3$ (blue, long dashed), and $N=2$ (green, short dashed). Parameters are $J=h=\Delta=1$ and $\beta=20$. For the collisional model we have fixed the interaction parameter $g\Delta t=\sqrt{Z_{a}/L}$ and for the Metropolis algorithm we have constructed the density matrix from averaging over $10^{6}$ runs.
In summary, the equivalence demonstrated here holds best at low temperatures and large system sizes of models that have relatively few degeneracies in their spectrum. The accuracy at large system sizes is particularly relevant, as these are precisely the regimes where the collisional model dynamics are most computationally expensive. Notably, while the calculation of the partition function $Z_{a}$ is computationally expensive, the partition function is not actually used in the Metropolis algorithm. It is only necessary to verify the equivalence. If there is a priori reason to believe a model has a well spaced spectrum with few degeneracies, and thus that $\sqrt{Z_{a}/L}$ is small, then $Z_{a}$ is not required to generate the dynamics from the Metropolis algorithm.
@@ V Concluding Remarks
In this work we have verified that both the collisional model framework and the Metropolis algorithm lead to thermalization of a many-body system when, in the case of the collisional model, the spectrum of the bath ancillae corresponds to each of the energy eigenstate transitions in the system. We have then demonstrated analytically that not only do both schemes produce thermalization, but that the time-dependent dynamics generated by both models are exactly equivalent when a condition relating the collisional model interaction strength and the ratio of bath ancilla partition function to number of possible energy eigenstate transitions is fulfilled.
The Metropolis algorithm is typically considered to be a purely phenomenological method of sampling possible system configurations, while collisional models are constructed to simulate underlying microscopic interactions that give rise to the system dynamics. These results demonstrate that, assuming the equivalence conditions are met, the Metropolis algorithm can, in fact, capture the same dynamics of the thermalization process as the collisional model. This could make the Metropolis model useful in situations where the transient dynamics leading up to thermalization are important, such as studying the performance of finite time quantum heat engines [54, 55, 56] or verifying stochastic thermodynamic behavior such as fluctuation theorems [57].
@@
Despite the similarities discussed here, there remain some important distinctions between the collisional model and the Metropolis algorithm. By microscopically modeling the system-bath interaction, the collisional model is significantly more general and can be applied with arbitrarily structured bath ancillae to study non-thermal steady states of open system dynamics.
In terms of computational cost, to simulate thermalizing dynamics both models require computing the energy eigenstates of the system, which is a computationally expensive process that scales exponentially with the system size. However, the Metropolis algorithm is advantaged in that it simply tracks individual eigenstate transitions and needs to know only about the Hilbert space of the system. The collisional model must take into account the much larger joint system-bath Hilbert space and perform repeated partial traces after each system-bath interaction. Thus the cost of the collisional model grows exponentially with the combined system plus bath ancilla dimensions, while the cost of the Metropolis scheme grows exponentially with the system size alone.
@@
For example, consider a system whose Hilbert space dimension is $d$. For modeling thermalizing dynamics using a bath whose spectrum has an excited state corresponding to each transition in the system, as we have done in our demonstration of thermalization in the XXZ model, the Hilbert space of each bath ancilla must be of dimension $\binom{d}{2}+1$. Thus the total Hilbert space that the collisional model must operate in is of dimension $\mathcal{O}(d^{3})$. Even in the most optimal scenario, when the spectra of the bath ancillae are identical to the spectrum of the system, the joint system bath Hilbert space will be of dimension $\mathcal{O}(d^{2})$. We must also keep in mind that the Metropolis scheme does incur an additional cost from the fact that the dynamics must be averaged over many repeated runs in order to construct the time-dependent density matrix, while the the collisional model directly models the density matrix evolution. However, each Metropolis experimental run still consists of only a series of products of $d$-dimensional operators.
@@
These results open several avenues for potential future work. Here we have shown an equivalence between the collisional model and Metropolis approaches in thermalizing dynamics for short collisional model interaction times and subject to the condition that the ratio of bath partition function to number of system energy eigenstate transitions is small. It may be possible to generalize these results to generic open system dynamics by modifying the distribution that the accept/reject probabilities are drawn in the Monte Carlo approach. Collisional models have seen extensive use in studying non-Markovian open system dynamics. It would be interesting to see if a similar equivalence could be found between non-Markovian collisional models and non-Markovian Monte Carlo schemes for quantum evolution [58, 59, 60].